PLEASE HELP ME SOLVE THS PROBLEM
IF THE TOTAL COST OF X APPLES IS B CENTS, WHAT IS A GENERAL FORMULA FOR THE COSTS, IN CENTS, OF Y APPLES?

Answer by stanbon(57361)

IF THE TOTAL COST OF X APPLES IS B CENTS, WHAT IS A GENERAL FORMULA FOR THE COSTS, IN CENTS, OF Y APPLES?
----------------------
Each apple cost x/b cents
So y apples will cost y(x/b) cents.
-------------------------------
You could also do this with a proportion:
Let "c" be the cost of y apples.
c/y = x/b
Multiply both sides by y to get:
c = y(x/b) cents
====================
Cheers,
Stan H.